November 1996  ICTCM IX (Reno)  12

Step 2 – Solution with Deployed Parachute

> IC2 := { x(td) = subs( t=td, X1 ),
>  v(td) = subs( t=td, V1 ) };

IC2 := \{v( td ) = − \frac{m g}{k1} + \frac{m g e^{(− \frac{td k1}{m} )}}{k1} ,

x( td ) = \frac{−m g td k1 + m^{2} g + x0 k1^{2} − m^{2} g e^{(− \frac{td k1}{m} )}}{k1^{2}}\}

> SOL2 := dsolve( subs(k=k2,SYS) union IC2, { x(t), v(t) } ):
> V2 := subs( SOL2, v(t) ):
> X2 := subs( SOL2, x(t) ):